Adjusted Gross Income
($1,000s)
Itemized Deductions
($1,000s)
22 9.6
27 9.6
32 10.1
48 11.1
65 15.5
85 19.7
120 25.5

a) Use the least squares method to develop the estimated regression equation that can be used to predict itemized deductions (in $1,000s) given the adjusted gross income (in $1,000s). (Round your numerical values to three decimal places.)

ŷ =

(b) Predict the reasonable level of total itemized deductions (in $1,000s) for a taxpayer with an adjusted gross income of $52,500. (Round your answer to two decimal places.)

$  thousand
Answer #1
X Y XY X^2 Y^2
22 9.6 211.2 484 92.16
27 9.6 259.2 729 92.16
32 10.1 323.2 1024 102.01
48 11.1 532.8 2304 123.21
65 15.5 1007.5 4225 240.25
85 19.7 1674.5 7225 388.09
120 25.5 3060 14400 650.25

PREDICTED Ý = a b X

From the above table and formula we get thevalue are as:

n 7
sum(XY) 7068.40
sum(X) 399.00
sum(Y) 101.10
sum(X^2) 30391.00
sum(Y^2) 1688.13
b 0.1707
a 4.7116

ycap = a+ bx

ycap = 4.712 + 0.171x

b)

when x = 52.5

ycap = 4.712 + 0.171 * 52.5

ycap = 13.69